OpenStudy (anonymous):

please help me solve this differential equation : g(t) =t^3/2

OpenStudy (anonymous):

Do you mean you want the derivative?

OpenStudy (anonymous):

Did you mean find the derivative?

OpenStudy (anonymous):

If so apply the power rule: y=t^3/2 y'=(3/2)t^(1/2)

OpenStudy (anonymous):

No , but using the laplace Transform table. Thanks

OpenStudy (anonymous):

Ohh. I got no clue then... Sorry.

OpenStudy (anonymous):

me either.